u buy 100 animals for $100, cows cost $10,pigs cost $1.75 and chickens cost $.50 how many can u buy
Answer by ikleyn(52778)

.

The Table below illustrates my attempts to get a solution by the "Trial and error" method.

Step by step I am approaching to the final solution.

It is presented in the last line.


 n     cows    pigs   chickens    money

 1      8      10        5        8*10 + 10*1.75 +  5*0.50 = 80 + 17.50 + 2.50 = 100

 2      6      20       10        6*10 + 20*1.75 + 10*0.50 = 60 + 35.00 + 5.00 = 100

 3      4      30       15        4*10 + 30*1.75 + 15*0.50 = 40 + 52.50 + 7.50 = 100

 4      2      40       20        2*10 + 40*1.75 + 20*0.50 = 20 + 70.00 +10.00 = 100

 5      0      50       25        0*10 + 50*1.75 + 25*0.50 =  0 + 87.50 +12.50 = 100

 6      0      40       60        0*10 + 40*1.75 + 60*0.50 =  0 + 70.00 +30.00 = 100     (*)


The last line gives the final solution to your problem.